Answer the following questions about the SO3
2- ion.
g. Draw the best Lewis structure that can be drawn with all single bonds.
h. Give the formal charge of each atom present.
i. Give the oxidation number of each atom present.
j. Re-draw the molecule with two single bonds and one double bond.
k. Give the formal charge of each atom in this second structure.
l. Give the oxidation number of each atom in this second structure.
Answer #1

done pair - no of eo con bond Ó I u FC= valence - po shell o 6-3-3 =0 = 6-6-1 - & US=0 @ so to